OpenStudy (anonymous):

(6n)/(n^2-4) * (n+2)/(n^2)

OpenStudy (anonymous):

\[\frac{ 6n }{n^2-4}*\frac{ n+2 }{ n^2 }\]

OpenStudy (anonymous):

Can you make factors for denominator of first term?

OpenStudy (anonymous):

\[a^2 - b^2 = (a+b)(a-b)\] This can help you to factorize it..